### v1

# MT-RewardTree: A Comprehensive Framework for Advancing LLM-Based Machine Translation via Reward Modeling

Abstract
Process reward models (PRMs) have shown success in complex reasoning tasks for large language models (LLMs). However, their application to machine translation (MT) remains underexplored due to the lack of systematic methodologies and evaluation benchmarks. To address this gap, we introduce MT-RewardTree, a comprehensive framework for constructing, evaluating, and deploying process reward models in MT. Unlike traditional vanilla preference pair construction, we propose a novel method for automatically generating token-level preference pairs using approximate Monte Carlo Tree Search (MCTS), which mitigates the prohibitive cost of human annotation for fine-grained steps. Then, we establish the first MT-specific reward model benchmark and provide a systematic comparison of different reward modeling architectures, revealing that token-level supervision effectively captures fine-grained preferences. Experimental results demonstrate that our MT-PRM-Qwen-2.5-3B achieves state-of-the-art performance in both token-level and sequence-level evaluation given the same input prefix. Furthermore, we showcase practical applications where PRMs enable test-time alignment for LLMs without additional alignment training and significantly improve performance in hypothesis ensembling. Our work provides valuable insights into the role of reward models in MT research. Our code and data are released in https://sabijun.github.io/MT_RewardTreePage.
MT-RewardTree: A Comprehensive Framework for Advancing LLM-Based Machine Translation via Reward Modeling
Zhaopeng Feng1♠ Jiahan Ren1♠  Jiayuan Su1♠  Jiamei Zheng1 Zhihang Tang1  Hongwei Wang1†  Zuozhu Liu1† 1Zhejiang University {zhaopeng.23,jiahan.24,jiayuan.23,jiamei.24,hongweiwang,zuozhuliu}@intl.zju.edu.cn
footnotetext: Equal contribution.footnotetext: Corresponding author.
Figure 1: Components of MT-RewardTree.
1 Introduction
The next-token prediction process in large language models (LLMs) is often modeled as a Markov Decision Process (MDP) and has achieved remarkable success across various domains, largely attributed to reinforcement learning (RL) and the scaling of test-time compute (Snell et al., 2024; Zeng et al., 2024; DeepSeek-AI et al., 2025; Team, 2025; Xiang et al., 2025). Reward models are central to these advancements. Outcome Reward Models (ORMs), which are designed to evaluate full responses, have been widely adopted; however, due to the sparsity of outcome rewards, ORMs often yield suboptimal performance and struggle with stability and efficiency during RL training (Lightman et al., 2024; Cao et al., 2024; Chan et al., 2024). In contrast, Process Reward Models (PRMs) evaluate intermediate steps to provide fine-grained guidance during both training and inference. PRMs have proven particularly effective in tasks such as mathematics and coding by guiding stepwise decision-making (Wang et al., 2024a; Chen et al., 2024; Luo et al., 2024; Qi et al., 2024; Guan et al., 2025).
Machine translation (MT) naturally aligns with token-level MDP frameworks, as each translation decision corresponds directly to token generation. Ramos et al. (2024) pioneered the use of the automatic metric xCOMET (Guerreiro et al., 2024) as a reward model to generate dynamic token-level rewards for RL training. Additionally, Zhao et al. (2024) and Wang et al. (2024b) demonstrated that scaling test-time compute can enhance LLM-based translation quality. However, these works do not fully address reward modeling, and there is still a lack of systematic methodologies for constructing and evaluating PRMs in MT, which has hindered progress relative to advancements in general-domain LLMs.
Developing effective PRMs is challenging. Although Lightman et al. (2024) demonstrate that process supervision with human annotators improves PRM performance in mathematical tasks, this methodology requires domain-expert annotators, resulting in prohibitive costs and practical limitations for translation tasks. Recently, some studies suggest that a PRM can be automatically learned during Direct Preference Optimization (DPO) training (Rafailov et al., 2024b, a; Yuan et al., 2024). However, existing vanilla preference pair datasets provide only sequence-to-sequence preference data, rather than token-level preferences, which raises concerns about their applicability for token-level alignment. Additionally, evaluating PRMs remains a significant challenge. In mathematical tasks, evaluation is often done using a Best-of-N (BoN) sampling strategy—selecting the highest-scored response from N candidates based on a PRM (Lightman et al., 2024; Wang et al., 2024c; Luo et al., 2024)—or by having the PRM identify errors or verify correctness in the steps (Zheng et al., 2024; Zhang et al., 2025). Since each step in mathematics has a deterministic answer, these methods do not directly translate to PRM evaluation in MT.
In this paper, we introduce MT-RewardTree, a comprehensive framework for constructing, evaluating, and deploying PRMs in machine translation. We propose an approximate Monte Carlo Tree Search (MCTS) method (Kocsis and Szepesvári, 2006; Silver et al., 2016) to generate the token-level preference pair dataset. This dataset is then split into a training set for reward model development and a benchmark for reward evaluation. We provide a systematic comparison of different reward modeling methods and test on both token-level and sequence-level performance. Furthermore, we demonstrate two practical applications of PRMs, including test-time alignment and hypothesis ensembling, offering valuable insights for future MT research. Our main contributions are as follows:
  • We construct the first reward model benchmark in MT and provide the first systematic comparison of reward modeling approaches for MT. Experimental results show that our MT-PRMs achieve competitive performance at both the token-level and sequence-level.
  • Our token-level preference pairs, generated through an approximate MCTS method, significantly outperform vanilla preference pairs in process reward model training.
  • We demonstrate that our MT-PRMs can facilitate test-time alignment for LLM-based MT without the need for additional alignment training. They can also be integrated for hypothesis ensembling to provide better translations, delivering performance parity with both MBR decoding and commercial LLM selection.
Figure 2: The construction process of token-level preference pairs. We utilize TowerInstruct-7B-v0.2 to generate candidate tokens. A token-level preference pair comprises two translations that share an identical prefix.
2 Background
2.1 Token-level Markov Decision Process
LLMs’ autoregressive generation can be naturally formulated as a Markov Decision Process, where each token generation is treated as an action. At each time step $t$, an action $a_{t}$ corresponds to the generation of a new token, and the state $\mathbf{s}_{t}$ is represented as the sequence of tokens generated up to that point. For tasks that do not involve interaction with an external environment—such as translation—the state is defined as
$\mathbf{s}_{t}=(x_{0},\dots,x_{L},y_{0},\dots,y_{t-1}),$
where $(x_{0},\dots,x_{L})$ represents the input prompt and $(y_{0},\dots,y_{t-1})$ is the sequence of generated tokens until time step $t-1$. The state transition function $f$ is deterministic and updates the state by concatenating the newly generated token:
$\mathbf{s}_{t+1}=f(\mathbf{s}_{t},a_{t})=\mathbf{s}_{t}\mid a_{t},$
with $\mid$ denoting concatenation.
Within this token-level MDP framework, the reward function $r(\mathbf{s}_{t},a_{t})$ is typically designed to provide feedback only at the terminal time step $T$, reflecting the overall correctness of the generated sequence or the successful completion of the task. To optimize the policy $\pi_{\theta}$ based on this reward, Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022) typically maximizes a KL-constrained objective:
$\mathbb{E}_{(s_{0},\ldots,s_{T})\sim\rho_{\pi}}\left[\sum_{t=0}^{T}\left(r(s_{t},a_{t})-\beta\log\frac{\pi(a_{t}\mid s_{t})}{\pi_{\text{ref}}(a_{t}\mid s_{t})}\right)\right],$
where $\pi_{\text{ref}}$ is a pre-trained reference policy, $\beta$ controls the strength of the KL penalty and $\rho_{\pi}$ denotes the trajectory distribution induced by policy $\pi$.
In practice, classical RLHF applies the reward solely at the terminal state. Specifically, the reward function used in Proximal Policy Optimization (PPO) (Schulman et al., 2017) is defined as:
$r(s_{t},a_{t})=\begin{cases}\beta\log\pi_{\text{ref}}(a_{t}\mid s_{t}),&\text{if }s_{t+1}\text{ is non-terminal},\\[2.84526pt] r(x,y)+\beta\log\pi_{\text{ref}}(a_{t}\mid s_{t}),&\text{if }s_{t+1}\text{ is terminal}.\end{cases}$
2.2 Reward Modeling in RLHF
Reward modeling is the cornerstone of RLHF, enabling LLMs to align their outputs with human preferences. In this section, we distinguish between typical (sequence-level) reward modeling and the more fine-grained token-level approach.
Sequence-level Reward Modeling. In classical RLHF, the reward function is learned from human feedback on prompt-response pairs $(\mathbf{x},\mathbf{y})$. The reward model is formulated as a contextual bandit, where a scalar reward is assigned only at the terminal state—i.e., once the full response sequence has been generated. This formulation, known as Outcome Reward Modeling, follows the Bradley-Terry (Bradley and Terry, 1952) preference model to define the probability of preferring one response over another:
$p^{*}(\mathbf{y}^{w}\succeq\mathbf{y}^{l})=\frac{\exp\left(r_{\phi}(\mathbf{x},\mathbf{y}^{w})\right)}{\exp\left(r_{\phi}(\mathbf{x},\mathbf{y}^{w})\right)+\exp\left(r_{\phi}(\mathbf{x},\mathbf{y}^{l})\right)}.$
To train the reward model $r_{\phi}$, we construct a preference dataset $\mathcal{D}$, where each prompt $x$ is paired with two candidate responses, $y$ and $y^{\prime}$. Human annotators or heuristics determine the preferred response $y_{w}$ and the rejected response $y_{l}$. The reward model is then optimized to maximize the likelihood of these human preferences:
$\max_{\phi}\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\sigma(r_{\phi}(x,y_{w})-r_{\phi}(x,y_{l}))\right],$
where $\sigma$ is the logistic function. By training $r_{\phi}$ in this manner, we ensure that the model assigns higher rewards to preferred responses, effectively capturing human-like quality judgments for sequence-level evaluation.
Token-level Reward Modeling. While the sequence-level approach treats the entire generated response as a single action, it fails to capture the fine-grained decision-making process inherent in token generation. Token-level reward modeling addresses this limitation by evaluating rewards at each token-generation step. This approach corresponds to a form of Process Reward Models. The cumulative reward for a trajectory $\tau$ is computed as the sum of per-token rewards, and the corresponding preference probability between two trajectories, $\tau^{w}$ and $\tau^{l}$, is given by:
$p^{*}(\tau^{w}\succeq\tau^{l})=\frac{\exp\left(\sum_{i=1}^{N}r(\mathbf{s}_{i}^{w},a_{i}^{w})\right)}{\exp\left(\sum_{i=1}^{N}r(\mathbf{s}_{i}^{w},a_{i}^{w})\right)+\exp\left(\sum_{i=1}^{M}r(\mathbf{s}_{i}^{l},a_{i}^{l})\right)}.$
Although token-level reward modeling offers finer-grained feedback, obtaining effective PRMs is more challenging to obtain and deploy compared to ORMs (Lightman et al., 2024; Cao et al., 2024).
3 MT-RewardTree
In this section, we introduce the components of the MT-RewardTree. We first describe how we construct token-level preference pairs using an MCTS-based method. Next, we review several approaches employed for reward modeling.
3.1 Constructing Token-level Preference Pairs
Prior studies have investigated translation preference pair construction (Xu et al., 2024; Agrawal et al., 2024; Feng et al., 2024a), yet a standardized token-level preference pair dataset for PRMs in MT remains absent. MQM (Freitag et al., 2021) datasets depend on manual error annotation, which is both cost-prohibitive and incapable of producing granular token-level preference pairs.
Drawing inspiration from MCTS, we propose a token-centric approach that quantifies token quality based on its potential to contribute to higher-quality translations. This method aligns with Monte Carlo-based PRMs construction techniques in mathematics, where step-wise quality is determined by its incremental contribution to deriving correct answers (Wang et al., 2024c; Guan et al., 2025).
The MCTS process consists of four main steps (depicted in Figure 2): Selection, Expansion, Simulation (Evaluation), and Back-propagation.
  1. 1.
    Selection: The first phase involves selecting a portion of the existing tree that is most promising for further expansion. Starting from the root node, a standard approach would traverse the tree down to a leaf using the PUCT algorithm (Rosin, 2011; Silver et al., 2017). Since our goal is to construct token-level preference pairs rather than achieving global optimality, we automatically select the existing prompt and previously generated tokens as the prefix $y_{<t}$.
  2. 2.
    Expansion: If the selected leaf node is not an EOS (end-of-sentence) token—i.e. if it is not a terminal state—the node is expanded by generating $k$ candidate children. This is achieved by decoding one additional step using the language model and selecting the top-$k$ tokens as the new children. We select the top-2 candidate tokens $a_{tj}$ (with $j\in\{1,2\}$) that have the highest logits. Preliminary experiments demonstrate that tokens outside of the top-2 yield significantly lower translation quality during the Simulation phase. These top-2 tokens, sharing the same prefix $y_{<t}$, form the basis for our token-level preference pair.
  3. 3.
    Simulation (Evaluation): From each expanded node $a$, we generate $n$ complete translation rollouts until an EOS token is reached. We then evaluate the quality (or groundedness) of the full translation sequence, denoted by $g(y,n)$. In our framework, we use COMETKiwi (Rei et al., 2022) to estimate the quality of all $n$ full rollouts. These scores are averaged and further assigned as the value of node $a$, i.e., $V(a)$.
  4. 4.
    Back-propagation: Since our objective is to construct token-level preference pairs, we compare the values $V(a_{t1})$ and $V(a_{t2})$ to determine which expanded token is superior. Finally, we retain the node with the higher $V$ value. This node, along with its corresponding prefix $y_{<t}$, is then used as the starting point in the next simulation cycle, beginning again at Step 1.
These four steps are repeated until the EOS token appears during the Selection phase. We retain one rollout from the superior token and one from the inferior to construct our token-level preference pair. We use COMETKiwi to guarantee the score gap lies between 0.04 and 0.4 to control the quality.
Model Sequence-level Token-level
Prefixed Arbitrary Prefixed
EN→XX XX→EN Avg. EN→XX XX→EN Avg. EN→XX XX→EN Avg.
Baselines
Skywork-Reward-LLaMA-3.1-8B 0.857 0.773 0.815 0.768 0.743 0.756 - - -
MT-Ranker-base 0.785 0.787 0.786 0.720 0.712 0.716 - - -
MT-Ranker-large 0.847 0.873 0.860 0.750 0.805 0.778 - - -
[…]
MT-PRM-LLaMA-3.2-3B 0.777 0.775 0.776 0.675 0.660 0.668 0.542 0.615 0.578
MT-PRM-Qwen-2.5-3B 0.867 0.858 0.863 0.622 0.633 0.628 0.637 0.685 0.660
Table 1: Accuracy results on RM-RewardBench. Skywork-Reward-LLaMA-3.1-8B is an advanced ORM for general domains, while MT-Ranker represents the SoTA non-metric reference-free translation quality estimation model.
Preference Pair Type Training Strategy Avg.
Token-level DPO 0.660
Vanilla DPO 0.574
Token-level KTO 0.644
Vanilla KTO 0.562
Table 2: Ablation study on the effect of training preference data and implicit reward training objectives. The backbone model is Qwen-2.5-3B-Instruct.
3.2 Implicit Process Reward Modeling
Unlike ORMs, which assign a single reward to the entire response, PRMs aim to assign rewards at a finer granularity, such as at each step or token. However, traditional PRMs training requires step-level annotations, which are costly to obtain. Recent studies (Rafailov et al., 2024a; Zhong et al., 2024) show that ORMs can be trained with implicit reward modeling, enabling PRMs to emerge naturally without the need for explicit step labels.
Consider an ORM where the reward is parameterized by the log-likelihood ratio of two causal language models:
$r_{\theta}(\mathbf{y}):=\beta\log\frac{\pi_{\theta}(\mathbf{y})}{\pi_{\text{ref}}(\mathbf{y})}$ (6)
where $\pi_{\theta}$ represents the trained model’s probability distribution, and $\pi_{\text{ref}}$ is a reference model. We define the cumulative reward up to step $t$ as:
$q_{\theta}^{t}(\mathbf{y}_{<t},y_{t}):=\sum_{i=1}^{t}\beta\log\frac{\pi_{\theta}(y_{i}|\mathbf{y}_{<i})}{\pi_{\text{ref}}(y_{i}|\mathbf{y}_{<i})}$ (7)
which serves as an exponential moving average of $r_{\theta}$ across steps. The expected process reward at step $t$ can then be expressed as:
$q_{\theta}^{t}(\mathbf{y}_{<t},y_{t})=\beta\log\mathbb{E}_{\pi_{\text{ref}}(\mathbf{y}|\mathbf{y}_{\leq t})}\left[e^{\frac{1}{\beta}r_{\theta}(\mathbf{y})}\right]$ (8)
This formulation shows that $q_{\theta}^{t}$ is an exact expectation of the outcome reward $r_{\theta}$ at step $t$, making it analogous to a Q-value in reinforcement learning.
By defining the process reward $r_{\theta}^{t}$ as the difference between successive Q-values:
$r_{\theta}^{t}:=q_{\theta}^{t}-q_{\theta}^{t-1}=\beta\log\frac{\pi_{\theta}(y_{t}|\mathbf{y}_{<t})}{\pi_{\text{ref}}(y_{t}|\mathbf{y}_{<t})}$ (9)
We see that PRMs can be derived directly from an ORM trained on response-level data, without requiring explicit step-wise labels. This insight suggests that training an ORM inherently leads to the learning of a Q-function, enabling step- or token-level reward modeling without requiring additional supervision. A typical example of this is DPO (Rafailov et al., 2024b), which optimizes the following objective:
$L_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$
This formulation shows that optimizing $\pi_{\theta}$ implicitly optimizes a reward model, as described in Eq. 4. Moreover, Yuan et al. (2024) demonstrated that this approach is agnostic to the specific training objective (i.e., not limited to DPO). It can be instantiated using various training objectives (e.g., KTO (Ethayarajh et al., 2024)), with the only modification being the substitution of $r_{\theta}(\mathbf{y})$ with $\beta\log\frac{\pi_{\theta}(\mathbf{y})}{\pi_{\text{ref}}(\mathbf{y})}$.
Moreover, our implicit PRMs can seamlessly be converted into ORMs using weighted implicit rewards:
$r_{\text{sequence}}(y_{1:T})=\sum_{k=0}^{T-1}w_{t}\log\frac{\pi_{\theta}(y_{t}|y_{<t})}{\pi_{\text{ref}}(y_{t}|y_{<t})}$ (11)
where the positional weights $w_{t}=\frac{1}{|y_{<t}|}$ are used to balance the contributions of each token.
4 Implementation Details
Datasets. We explore four languages—English (EN), German (DE), Chinese (ZH), and Russian (RU)—and six translation directions: EN→XX and XX→EN. Our raw corpus consists of test sets from WMT17 to WMT20, supplemented with development and test sets from the Flores (Costa-jussà et al., 2022). We use the TowerInstruct-7B-v0.2111https://huggingface.co/Unbabel/TowerInstruct-7B-v0.2 model with a temperature of 0.95 and apply the MCTS-based approach described earlier. During the Simulation step, we sample three candidate hypotheses for each node.
The token-level preference pair dataset is divided into two subsets: 200 pairs per translation direction are sampled as the token-level MT-RewardBench (Prefixed), which also serves as the sequence-level MT-RewardBench (Prefixed). The remaining pairs are assigned as token-level training data. Additionally, we sample two translations for each source sentence in the token-level datasets, ensuring they do not strictly share the same prefixed text (namely the vanilla preference pair). These pairs form the sequence-level arbitrary preference pair set. Detailed statistics can be found in Table 5.
Training Details. We take LLaMA-3.2-3B-Instruct [^2] https://huggingface.co/meta-LLaMA/LLaMA-3.2-3B-Instruct and Qwen-2.5-3B-Instruct [^3] https://huggingface.co/Qwen/Qwen2.5-3B-Instruct as the backbone models for training. For the DPO training, the higher-scored sentence is designated as the chosen response, while the lower-scored sentence is labeled as the rejected response. For the KTO training, the higher-scored sentence is treated as the positive sample, and the lower-scored sentence as the negative sample. We set $\beta$ as 0.1.
Reward Evaluation. We evaluate reward models by framing the task as a classification problem, similar to prior work on reward model benchmarks in the general domain (Lambert et al., 2024; Liu et al., 2024). For sequence-level evaluation, given a tuple $(x,y_{c},y_{r})$, where $x$ is the prompt, $y_{c}$ is the chosen response, and $y_{r}$ is the rejected response, the reward model predicts whether $y_{c}$ is better than $y_{r}$. If the reward model assigns a higher reward to $y_{c}$ than to $y_{r}$, the prediction is correct; otherwise, it is incorrect. We use accuracy as the evaluation metric, computed as follows:
$\text{Accuracy}=\frac{1}{|D|}\sum_{(x,y_{c},y_{r})\in D}I[R_{\theta}(x,y_{c})>R_{\theta}(x,y_{r})]$
where $I(\cdot)$ is the indicator function, and $D$ denotes the evaluation dataset.
For token-level evaluation, we use tuples of the form $(x,y_{<t}+a_{c},y_{<t}+a_{r})$, where $y_{<t}$ is the generated tokens before, $a_{c}$ is the next chosen token, and $a_{r}$ is the rejected token. Similarly, we compute accuracy as the evaluation score: if the PRM assigns a higher reward to $a_{c}$ than to $a_{r}$, the prediction is correct; otherwise, it is incorrect.
5 Evaluation Results
Token-level Performance. From Table 3.1, we can observe that our MT-PRM-LLaMA-3.2-3B and MT-PRM-Qwen-2.5-3B models achieved accuracies of 0.578 and 0.66 respectively on the token-level MT-RewardBench. As shown in Table 2, we systematically compare models trained with vanilla sequence-level preference pairs versus our token-level preference pairs, while evaluating both DPO and KTO training objectives. The results demonstrate that token-level preference pairs significantly improve discrimination accuracy: implicit PRMs trained with token-level preference pairs outperform vanilla sequence-level baselines by +8.6% (DPO) and +11.5% (KTO). This performance gap highlights the critical advantage of token-level preference pairs in helping capture fine-grained translation quality distinctions.
Sequence-level Performance. We also convert our PRMs to sequence-level scoring through weighted DPO rewards (as shown in Eq. 11). We can observe that our MT-PRM-Qwen-2.5-3B achieves the highest performance among all models in the Prefixed set, with an average score of 0.863, outperforming both Skywork-Reward-LLaMA-3.1-8B  [^4] https://huggingface.co/Skywork/Skywork-Reward-Llama-3.1-8B and the MT-Ranker (Moosa et al., 2024) variants. This demonstrates the effectiveness of our token-level supervision framework even when adapted to sequence-level scoring. However, in the Arbitrary set, we observe a relative drop in performance (average accuracy between 0.628 and 0.668). We hypothesize this drop stems from the inclusion of noisy partial translations in the Arbitrary set, which disrupts the causal dependency structure crucial for token-level rewards, and the limited compositional reasoning capabilities of current PRM architectures when handling fragmented inputs, a challenge well-documented in partial sequence evaluation (Wang et al., 2024a; Yuan et al., 2024).
Strategy Model ZH-EN EN-ZH Avg.
XCOMET COMETKiwi XCOMET COMETKiwi
MBR BLEU 0.8243 0.7842 0.8106 0.7768 0.7990
BERTScore 0.8277 0.7874 0.8227 0.7837 0.8054
COMET 0.8371 0.7932 0.8334 0.7949 0.8147
[…]
MT-PRM-Qwen2.5-3B 0.8250 0.7866 0.8153 0.7841 0.8028
MT-PRM-LLaMA3.2-3B 0.8282 0.7908 0.8225 0.7918 0.8083
Table 3: Automatic evaluation metrics for different ensembling strategies across WMT 23 ZH-EN and EN-ZH.
Source {CJK*}UTF8gkai 换油的师傅说油品清亮,确实是好油。
[…]
Table 4: Case study of test-time alignment.
6 Practical Insights
6.1 Test-time Alignment
Task setup. Test-time alignment, also known as decoding-time alignment (Huang et al., 2024), refers to the process of adjusting an LM’s output during inference to better align with human preferences, without additional training or fine-tuning. Its application in MT remains underexplored.
In the context of MT, given the prior context $s_{<t}$ and timestamp $t$, we define the reward-guided scoring function for a candidate token $a$ as:
$\ s(a,s_{<t})=\text{LM}(a\mid s_{<t})+w\cdot P(r([s_{<t},a]))$ (13)
where $\text{LM}(a\mid s_{<t})$ represents the LM’s predicted probability for token $a$ given the preceding context $s_{<t}$. $r([s_{<t},a])$ denotes the reward signal for token $a$, conditioned on the prior context $s_{<t}$. The softmax function is applied over the reward signal $r([s_{<t},a])$, computed over the top $k$ candidate tokens (with $k$ being a window size), normalizing the reward value, which we label as $P(r([s_{<t}$, a]). The scaling factor $w$ adjusts the relative weight of the reward signal, allowing it to contribute effectively without overpowering the LM’s probability. Compared to standard decoding strategies, this approach offers a more refined scoring function, as it encourages the generated text to: 1) Maintain semantic coherence and relevance with the prior context, and 2) Align more closely with reward-based criteria and human preferences. Test-time alignment also substantially reduces the need for the extensive resources typically required for LM alignment training.
Results. We use Qwen2.5-14B-Instruct [^5] https://huggingface.co/Qwen/Qwen2.5-14B-Instruct for generating tokens and leverage MT-PRM-LLaMA-3.2-3B and MT-PRM-Qwen-2.5-3B as the models for providing token-level rewards. We randomly sample 500 cases from the WMT 2023 testset. As shown in Figure 3, the reward-guided decoding methods outperform the standard greedy decoding in both EN-RU and ZH-EN translation tasks, evaluated by the COMET (Rei et al., 2020), COMETKiwi (Rei et al., 2022), and XCOMET-XL (Guerreiro et al., 2024) metrics. For instance, using the XCOMET-XL metric, LLaMA PRM and Qwen PRM outperform the standard greedy decoding by 17.5% and 17.9% in the EN-RU task respectively. Additionally, Qwen PRM slightly outperforms LLaMA PRM in both translation tasks and across all metrics, which aligns with the results in Table 3.1, where Qwen PRM achieves better token-level reward performance. These findings highlight the effectiveness of reward-guided decoding strategies in improving MT outcomes.
Case Study. The case in Table 5 compares a ZH-EN translation using Greedy Decoding (GD), GD with LLaMA PRM, and GD with Qwen PRM. The standard GD translation, “The mechanic said the oil was clear, indeed good oil” , conveying the basic meaning but lacking the important "changing oil" context. Both reward-guided decoding methods improve the translation to a more complete, high-quality version, achieving a COMETKiwi score improvement of 4.97%.
Figure 3: Results of test-time alignment across WMT 23 ZH-EN and EN-RU. MT-PRMs with less parameters can assist in aligning Qwen-2.5-14B-Instruct.
6.2 Hypothesis Ensembling
Task Setup. Ensembling is widely recognized for its ability to combine multiple complementary models to improve performance in machine learning (Hansen and Salamon, 1990; Ting and Witten, 1997). Farinhas et al. (2023) explore generating multiple translation hypotheses and ensembling them to produce a single higher-quality translation. In this work, we investigate two complementary ensembling approaches for MT: ranking-based selection and Minimum Bayes Risk decoding.
In ranking-based ensembling, the optimal output is selected via:
$\hat{y}_{\mathrm{ranking}}:=\mathbf{argmax}_{y\in\mathcal{\bar{Y}}}~{}f(y)$ (14)
where $f$ represents either: (1) Strong LLM as a judge (e.g., Gemini-2.0-Flash [^6] https://ai.google.dev/gemini-api/docs/models) or (2) Our PRM, converted to sequence-level scoring through weighted DPO rewards (as shown in Eq. 11).
For MBR decoding, we maximize the expected utility over candidate references:
$\hat{y}_{\mathrm{mbr}}:=\mathbf{argmax}_{y\in\bar{\mathcal{Y}}}\frac{1}{M}\sum_{j=1}^{M}u(y^{(j)},y)$ (15)
where $u(\cdot)$ can be instantiated using traditional metrics such as BLEU (Papineni et al., 2002), BERTScore (Zhang et al., 2019), and the reference-based metric COMET (Rei et al., 2020) for upper-bound analysis. This dual-strategy framework allows for a comprehensive evaluation of the ensembling potential of reward models.
Results. We evaluate on 500 cases sampled from the WMT 2023 dataset, using the TowerInstruct-7B-v0.2 model with nucleus sampling to generate 8 candidate translations for each case. As shown in Table 5, our MT-PRM-LLaMA-3.2-3B outperforms MBR decoding, based on BLEU and BERTScore, by 0.93% and 0.27%, respectively. It even surpasses the commercial LLM Gemini-2.0-Flash by 0.51%. When comparing the performance of different PRMs, we find that LLaMA outperforms Qwen, which is consistent with their relative performance in the MT-RewardBench sequence-level evaluation for the Arbitrary set. This further validates the practical utility of MT-RewardBench.
7 Related Work
Token-Level Feedback Mechanisms. Fine-grained feedback has been recognized for its ability to help models capture potential errors more precisely (Lightman et al., 2024). In the context of mathematical reasoning, process supervision using Monte Carlo methods has shown significant promise (Wang et al., 2024c; Qi et al., 2024; Guan et al., 2025). Furthermore, developments in general-domain have demonstrated that DPO can implicitly learn token-level rewards through policy optimization, a process referred to implicit reward learning (Rafailov et al., 2024a; Wang et al., 2024a; Yuan et al., 2024). Despite these advancements, these approaches have yet to be tested in the context of MT. The translation community has long acknowledged the value of granular feedback, with early attempts relying on binary error markings from human annotations (Kreutzer et al., 2020), reference-based heuristics (Petrushkov et al., 2018), or LLM (Feng et al., 2024b).
Alignment Paradigms in Machine Translation. Alignment techniques in neural machine translation have evolved from Minimum Risk Training (Shen et al., 2015) to more sophisticated reinforcement learning approaches (Dang et al., 2024). While PPO-based RLHF has achieved success in general-domain alignment, its application to MT presents unique challenges, particularly due to the need for fine-grained quality signals rather than the bandit reward. Recent works like He et al. (2024) and Xu et al. (2024) have investigated the use of automatic metrics to select better translations or construct preference pairs to improve the LLM, while Zhao et al. (2024) explored scaling test-time compute to further enhance translation performance. Recently, Ramos et al. (2024) pioneered the use of xCOMET as a dynamic reward signal during RL training. However, these methods remain limited to sequence-level guidance or binary approximations of the reward process, failing to provide the fine-grained token-level feedback required for more accurate translation alignment. MT-RewardTree provides a new perspective by introducing a more granular, token-level reward modeling framework.
8 Conclusion
In this work, we propose MT-RewardTree, a comprehensive framework for constructing, evaluating, and deploying process reward models in machine translation. Our framework leverages an automatic token-level preference pair generation approach inspired by approximate Monte Carlo Tree Search, effectively addressing the challenge of large-scale fine-grained supervision annotation. Extensive experiments on both sequence-level and token-level benchmarks demonstrate that our MT-PRM achieves advanced performance in reward modeling in MT, surpassing traditional sequence-level preference pairs. Our exploration of test-time alignment and hypothesis ensembling provide valuable insights for the application of reward models.
Limitations
Although we have developed the first comprehensive framework for process reward models in the field of machine translation, several important challenges remain to be addressed. Our work primarily focuses on synthesizing token-level data to leverage its fine-grained benefits. However, methods like Token-level DPO, RTO which optimize training algorithms, also show promise in further improving PRM performance. Additionally, our current framework includes only a limited set of high-resource languages, and expanding to multilingual settings, especially for low-resource languages, is a crucial direction for future work. While we have demonstrated the potential applications of reward models in test-time alignment and hypothesis ensembling, their integration into reinforcement learning training remains an important area for exploration.